# Measuring Human and AI Values based on Generative Psychometrics with Large Language Models

### v1

@@ 4.2 Case Study
We exemplify the advantage of GPV over prior data-driven tools such as PVD in Fig. 3. Some values, while not explicitly mentioned in PVD-designed lexicons, are implied within the text. For example, in Schwartz’s theory, Achievement is defined as "the personal pursuit of success, demonstrating competence according to social standards." In this context, "the teacher’s praise" and "performing well in an exam" both embody the "success" element of achievement. Although the text does not directly reference Achievement or Achievement-related lexicons, the author’s expression of joy and aspiration for these outcomes reflects this value. While GPV effectively captures this aspect, PVD does not.
Some PVD-designed lexicons fail to reflect their intended values or align with the measurement subject in context. For instance, words like "friendly" and "goal" describe the author’s deskmate, and "money" refers to something the author found, but none of these terms convey the author’s values. GPV effectively avoids such misinterpretation.
@@ 5 GPV for Large Language Models
We evaluate 17 LLMs across 4 value systems using three measurement tools: self-report questionnaires [34], ValueBench [63], and GPV. Unless otherwise specified, we use LLM-generated value-eliciting questions for GPV to ensure a comprehensive and thorough measurement of each value. The detailed experimental setup is described in § D.1.
@@ 5.1 Comparative Analysis of Construct Validity
Figure 4: Correlation between Schwartz values when using different measurement tools.
Using the measurement results from 17 LLMs as data points, we compute the correlation between each pair of values in Schwartz’s value system. The correlation results are visualized in a heatmap for each measurement tool in Fig. 4. The heatmap reveals the superior construct validity of GPV, as its measurement results align more closely with the theoretical structure (Fig. 2(a)). Specifically, values that are adjacent in the theoretical circumplex structure exhibit positive correlations, while those that are theoretically distant show negative correlations.
In contrast, most of the value pairs generated by prior tools [34, 63] exhibit positive correlations, contrary to theoretical expectations. This discrepancy indicates a strong susceptibility to response biases in these tools, wherein certain LLMs generally tend to assign higher scores in self-report measures or respond more supportively in ValueBench assessments than others. Such biases obscure the genuine value orientations of the LLMs.
Besides Schwartz’s value system, we also evaluate the construct validity by relating the values of different value theories that are theoretically positively correlated. Results in Table 4 indicate the superior construct validity of GPV.
In summary, evaluations within and across value theories indicate superior construct validity of GPV over prior tools [34, 63] that are prone to response bias.
@@ 5.2 Comparative Analysis of Value Representation Utility
The utility of human value measurements lies in their predictive power for human behavior [83]. In the context of LLMs, many related studies are motivated by the value alignment for their safe deployment [35, 97]. However, few studies have connected LLM values with their safety. In this section, we evaluate the value representation utility of different measurement tools in terms of their predictive power for LLM safety scores.
Here, we use the safety scores of 17 LLMs from SALAD-Bench [43] as the ground truth and randomly sample 100 prompts from Salad-Data [43] for GPV measurement. We train a linear probing classifier to predict the relative safety of each pair of LLMs, using the value measurement results as features. We perform its training 30 times for each measurement tool with randomly sampled data splits to ensure statistically meaningful results. Full experimental details are given in § D.3.
Using different value systems as features leads to different results. We present the best classification accuracy of different measurement tools in Table 5. The results show that GPV is more predictive of LLM safety scores than prior tools. It suggests that GPV values can be an interpretable and actionable proxy for LLM safety under specific context [67].
@@ 5.3 Discussions
We discuss that the superior construct validity can be attributed to the knowledge embedded within ValueLlama and its ability to inject such knowledge into the measurement process. During pertaining and our fine-tuning, ValueLlama learns the correlations between different values. Therefore, when measuring LLM values, ValueLlama can leverage such knowledge to generate more coherent value representations, ensuring the construct validity of the measurement results. The superior value representation utility of GPV may be attributed to the context-specific value measurements. Values are relatively stable traits for humans. However, we may not treat LLMs as monolithic entities with consistent values like humans, which emphasizes the need for context-specific evaluations [67]. GPV, for the first time, enables such context-specific measurements. Overall, compared to prior tools, using GPV for LLM value measurements (1) mitigates response bias and yields more theoretically valid results; (2) is more practically relevant for measuring LLM values based on their scalable and free-form responses; and (3) enables context-specific measurements.
@@
The current studies are limited to the evaluation of LLMs in English. Since the used languages are shown to affect the values of LLMs [17], future research should consider multi-lingual measurements. Additionally, future investigations should explore the spectrum of values an LLM can exhibit, examining how different profiling prompts affect this spectrum. Nevertheless, as Röttger et al. [67] suggest, current alignment algorithms establish default model positions and behaviors, making it meaningful to evaluate the values and opinions reflected in these defaults. It is plausible that future models, as a result of more comprehensive alignment, will display more consistent value profiles and reduced variability in their responses to different prompts.
@@ Appendix A ValueLlama
Our training dataset comprises ValueBench [63] and ValuePrism [86]. ValueBench compiles 453 value dimensions and two thousand items from 44 established psychometric inventories. Item-value pairs in ValueBench are labeled as relevant, while an equivalent number of irrelevant pairs are randomly sampled. Valence labels are assigned according to agreement labels in ValueBench. 50 values and their corresponding items are excluded from the training dataset for evaluation purposes. ValuePrism consists of contextualized values generated by LLMs, linked to real-life situations written by humans. The relevance and valence labels in ValuePrism are also generated by LLMs, with their quality being validated by human annotators.
Prompt Templates.
In LABEL:prompt:_generation, we collect the used prompt templates for generating relevance and valence.
@@ B.1 Parsing Text into Perceptions
We employ semchunk [16] to recursively divides text into chunks of a specified size. It splits text using meaningfully ordered splitters (newlines, punctuations, etc.) and continues splitting until all chunks are within the desired size limit (250 tokens in this work). It then merges any smaller chunks until they reach the target size. Splitters that are not whitespace are reattached to the end of chunks if this does not exceed the size limit; otherwise, they are treated as separate chunks.
@@
LABEL:prompt:_parsing presents the prompt used for parsing perceptions.
@@ Appendix C GPV for Humans
C.1 Stability Analysis
With results presented in Table 6, we evaluate the stability of GPV, i.e., the alignment between perception-level and aggregated measurement results for humans. The results show that the perception-level measurement results are generally consistent with the individual-level ones, indicating desirable stability. Since values are defined as desirable end states [69], the perception-level measurements are more likely to support values than oppose them.
@@ C.1 Stability Analysis
C.2 Construct Validity
For each blog dataset, we first measure Schwartz Values using GPV, resulting in a 10-dimensional output for each data entry. We analyze the results from 791 entries by calculating cosine similarity, which produces a 10x10 matrix representing the similarity between the values. Individuals with unmeasured dimensions are excluded when calculating cosine similarity between each pair of two values. To ensure that higher similarity corresponds to smaller distances in the subsequent MDS analysis, each matrix element $\alpha$ is transformed as $1-\alpha$. Finally, MDS analysis is performed on the distance matrix to obtain the corresponding results.
C.3 Concurrent Validity
@@ D.1 Experimental Details
We measure values for 17 LLMs: internlm2-chat-7b (inte2) [18], internlm-chat-7b (inte) [18], Llama-2-7b-chat-hf (Lla2) [89], gemma-2b (ge2b) [88], gemma-7b (ge7b) [88], Qwen1.5-4B-Chat (Qw4B) [7], Qwen1.5-14B-Chat (Qw14B) [7], Qwen1.5-72B-Chat (Qw72B) [7], Qwen1.5-7B-Chat (Qw7B) [7], Qwen1.5-0.5B-Chat (Qw0.5B) [7], Qwen1.5-1.8B-Chat (Qw1.8B) [7], gpt-4-turbo (gpt4) [2], gpt-3.5-turbo (gpt3.5) [2], Yi-6B-Chat (Yi6B) [99], Mistral-7B-Instruct-v0.1 (Mis0.1) [36], Mistral-7B-Instruct-v0.2 (Mis0.2) [36], vicuna-7b-v1.5-16k (vic7b) [101]. All LLMs are set to greedy decoding mode. If applicable, the system prompts are set to "You are a helpful assistant." by default [103].
@@
Using LABEL:prompt:_item_generation, we ask GPT-4o to generate five open-ended, value-eliciting questions designed to evaluate each value dimension. In this manner, we ensure a comprehensive and sufficient measurement of each involved value. To tailor these questions effectively, we provide several guidelines, which may overlap in some aspects: (1) Avoid personal experience and emotion-based questions: Since an LLM doesn’t have personal experiences or emotions, questions should not assume it does. (2) Use value dilemmas and scenarios: Instead of asking direct questions that may lead to generic responses, use scenarios or dilemmas to reveal the LLM’s values. (3) Present ethical dilemmas: Include scenarios with no clear right or wrong answer to understand the model’s approach to complex moral issues. (4) Involve conflicting values: Create situations requiring a choice between conflicting values to reveal the model’s value prioritization. (5) Justification: After each scenario, ask the model to justify its decision to understand the reasoning behind it. We exemplify the generated questions below:
@@ D.2 Stability Analysis
Table 8 gathers the results for analyzing stability of GPV, i.e., the alignment between perception-level and aggregated measurement results for LLMs. They demonstrate that the perception-level results are generally consistent with the aggregated LLM-level results, indicating desirable stability.
@@
D.3 Comparative Analysis of Value Representation Utility
@@
We add a linear layer on top of the value measurement results to predict the relative safety scores of LLMs. The linear layer maps the value measurement results to the scalar safety score. LLMs are paired as data points; for each pair of LLMs, we concatenate their scores output by the linear layer as the classification logits. We use the cross-entropy loss to train the linear layer.
@@
To ensure a fair comparison, we normalize the value measurement results of different tools to $[-1,1]$ before feeding them into the linear layer. The missing value measurement results (due to refusal to answer or invalid responses in self-reports) are set to 0. We use the Adam optimizer with a learning rate of 0.001 to train the linear layer. We train the linear layer for 1000 epochs with the batch size equal to the dataset size (36 pairs). The checkpoint with the best validation accuracy is selected as the final model.
@@
Since VSM is the most predictive value system for LLM safety scores, we further investigate the contributions of different values in VSM to the predicted LLM safety scores. The linear probing parameter associated with each value, averaged over 30 runs of different data splits, are presented in Table 10. We find that Long Term Orientation, Indulgence, and Uncertainty Avoidance positively contribute to the predicted safety scores, while Masculinity, Power Distance, and Individualism negatively contribute. These findings provide insights into future value alignment strategies for LLM safety. Instead of relying human preferences, we may leverage the basic values of LLMs to guide their behavior towards safety, which can be more transparent, adaptable, and interpretable [97].
@@
Table 10: The associated parameters in the linear probing model for different values.
D.4 Value Orientations of LLMs

### v2

@@ 4.2 Case Study
We exemplify the advantage of GPV over prior data-driven tools such as PVD in Fig. 3. Some values, while not explicitly mentioned in PVD-designed lexicons, are implied within the text. For example, in Schwartz’s theory, Achievement is defined as “the personal pursuit of success, demonstrating competence according to social standards." In this context, “the teacher’s praise" and “performing well in an exam" both embody the “success" element of achievement. Although the text does not directly reference Achievement or Achievement-related lexicons, the author’s expression of joy and aspiration for these outcomes reflects this value. While GPV effectively captures this aspect, PVD does not.
Some PVD-designed lexicons fail to align with the measurement subject or reflect their intended values. For instance, “friendly" and “goal" target the author’s deskmate; picking up “money" does not indicate the author’s own values of Power. GPV effectively avoids such misinterpretation.
@@ 5 GPV for Large Language Models
We evaluate 17 LLMs across 4 value systems using 3 measurement tools: self-report questionnaires [35], ValueBench [65], and GPV. Unless otherwise specified, we use LLM-generated value-eliciting questions for GPV to ensure a comprehensive and thorough measurement of each value. The detailed experimental setup is described in § D.1.
@@ 5.1 Comparative Analysis of Construct Validity
Figure 4: Correlations between Schwartz values when using different measurement tools.
Using the measurement results from 17 LLMs as data points, we compute correlations between Schwartz’s values. The results are visualized in a heatmap for each measurement tool in Fig. 4. The heatmap reveals the superior construct validity of GPV, as its measurement results align more closely with the theoretical structure (Fig. 2(a)). Specifically, values that are adjacent in the theoretical circumplex structure exhibit positive correlations, while those that are theoretically distant show negative correlations.
In contrast, prior tools obtain almost all-positive correlations, contrary to theoretical expectations. This discrepancy indicates their strong susceptibility to response biases, wherein certain LLMs generally tend to assign higher scores in self-report or respond more supportively in ValueBench. Such biases obscure the genuine value orientations of the LLMs. Even when centering the measurement results of prior tools (see § D.3), the correlation results remain inconsistent with the theoretical structure. This finding aligns with recent studies revealing the unreliability of LLMs as survey respondents [23, 69].
Besides Schwartz’s value system, we also evaluate the construct validity by relating the values of different value theories that are theoretically positively correlated. Results in Table 4 indicate the superior construct validity of GPV; i.e., for the theoretically positively correlated values, measuring with GPV also yields higher correlations.
In summary, evaluations within and across value theories indicate superior construct validity of GPV over prior tools that are prone to response bias.
@@ 5.2 Comparative Analysis of Value Representation Utility
The utility of human value measurements lies in their predictive power for human behavior [83]. In the context of LLMs, many related studies are motivated by value alignment for safe LLM deployment [37, 98]. However, few studies have connected LLM values with their safety. In this section, we evaluate the value representation utility of different measurement tools in terms of their predictive power for LLM safety scores.
Here, we use the safety scores of 17 LLMs from SALAD-Bench [45] as ground truth and randomly sample 100 prompts from Salad-Data [45] for GPV measurement. We follow the standard linear probing protocol and train a linear classifier to predict the relative safety of LLMs, using the value measurement results as features. We perform its training 30 times for each measurement tool with randomly sampled data splits to ensure statistically meaningful results. Full experimental details are given in § D.4.
Using values from different value theories as features leads to different results. We present the best classification accuracy of different measurement tools in Table 5. The results indicate that GPV is more predictive of LLM safety scores than prior tools. It suggests that GPV values can be an interpretable and actionable proxy for LLM safety under specific context [69].
@@ 5.3 Discussion
We discuss that the superior construct validity may be attributed to the encoded knowledge. During pertaining and our fine-tuning, ValueLlama learns the correlations between different values, which is exploited to generate more coherent and valid measurements. In addition, measuring the free-form LLM responses is more reliable than prompting with forced-choice questions [23]. The superior value representation utility of GPV may be attributed to the context-specific value measurements. Unlike humans, who exhibit stable values, LLMs may not be treated as monolithic entities, highlighting the importance of context-specific measurement [69]. GPV, for the first time, enables reliable context-specific measurements. Overall, compared to prior tools, using GPV for LLM value measurements 1) mitigates response bias and yields more theoretically valid results; 2) is more practically relevant due to measuring scalable and free-form LLM responses; and 3) enables context-specific measurements.
@@
The current studies are limited to evaluating LLMs in English. Since the used languages are shown to affect LLM values [16], future research should consider multi-lingual measurements. Additionally, future investigations should explore the spectrum of values an LLM can exhibit, examining the effects of different profiling prompts. Though LLM values may be steerable, current alignment algorithms establish default model positions and behaviors, making it still meaningful to evaluate the values and opinions reflected in these defaults [69].
@@ 6 Conclusion
Ethical Statement
Measuring values with GPV may involve biases encoded in LLMs, during perception-level measurement and perception parsing. Currently, GPV is intended for research purposes only, and researchers should exercise caution when applying it to content with subjective or controversial interpretations.
For the perception-level measurement, we fine-tuned our model using established psychological inventories and synthetic data validated across cultures, aiming to reduce measurement bias. In the three-class valence classification task, the model is trained to provide neutral predictions when additional context is needed, thereby minimizing the risk of bias. Nevertheless, achieving unbiased measurement requires further investigation.
The parsing results in this study are considered high-quality by our annotators. However, since the annotators share a similar demographic background, their evaluations may lack a comprehensive and diverse perspective. Additionally, the blog data analyzed in this work primarily focuses on general, everyday topics and rarely involves controversial issues. Addressing potential biases in parsing remains an open area for future research.
Acknowledgements
This work is supported by the National Natural Science Foundation of China (Grant No. 62276006); Wuhan East Lake High-Tech Development Zone National Comprehensive Experimental Base for Governance of Intelligent Society; and the Fundamental Research Funds for the Central Universities.
Measuring Human and AI Values Based on Generative Psychometrics with Large Language Models
@@ Appendix A ValueLlama
Our training dataset comprises ValueBench [65] and ValuePrism [87]. ValueBench compiles 453 value dimensions and two thousand items from 44 established psychometric inventories. Item-value pairs in ValueBench are labeled as relevant, while an equivalent number of irrelevant pairs are randomly sampled. Valence labels are assigned according to agreement labels in ValueBench. 50 values and their corresponding items are excluded from the training dataset for held-out evaluation purposes. ValuePrism consists of contextualized values generated by LLMs, linked to real-life situations written by humans. The relevance and valence labels in ValuePrism are also generated by LLMs, with their quality being validated by human annotators.
Prompting.
We collect the used prompt templates in LABEL:prompt:_generation for generating relevance and valence. Admittedly, better prompting or implementing RAG for more powerful models may be valid alternatives to fine-tuning, our open-source, small fine-tuned model aims to be more accessible and scalable for wide adoption and large-scale studies. We leave model compression for future work.
@@ B.1 Parsing Text into Perceptions
We employ semchunk [15] to recursively divide text into chunks of a specified size. This work uses a chunk size of 250 tokens, which is a relatively small chunk size for the sake of higher-quality parsing.
@@
LABEL:prompt:_parsing is used for parsing perceptions.
@@ Appendix C GPV for Humans
C.1 Data Filtering
We filter out low-quality blogs from the Blog Authorship Corpus [77], which originally contains 9660 blogs, using the following criteria: gender field is not empty, word count > 1000, and the text does not contain "http", "urlLink", ":)", "*", "=)", "&nbsp", or "<U". After filtering, we obtain 791 blogs for further analysis.
C.2 Stability Analysis
With results presented in Table 6, we evaluate the stability of GPV, i.e., the consistency between perception-level and aggregated measurements. The results show that the perception-level measurement results are generally consistent with the individual-level ones, indicating desirable stability. Since values are defined as desirable end states [71], the perception-level measurements are more likely to support values than oppose them.
@@ C.2 Stability Analysis
C.3 Construct Validity
For each blog dataset, we first measure Schwartz Values using GPV, resulting in a 10-dimensional output for each data entry. We analyze the results from 791 entries by calculating cosine similarity, which produces a 10x10 matrix representing the similarity between the values. Individuals with unmeasured dimensions are excluded when calculating cosine similarity between each pair of two values. To ensure that higher similarity corresponds to smaller distances in the MDS analysis, each matrix element $\alpha$ is transformed as $1-\alpha$. Finally, MDS analysis is performed on the distance matrix to obtain the corresponding results.
C.4 Concurrent Validity
@@ D.1 Experimental Details
We measure values for 17 LLMs: internlm2-chat-7b (inte2) [17], internlm-chat-7b (inte) [17], Llama-2-7b-chat-hf (Lla2) [90], gemma-2b (ge2b) [89], gemma-7b (ge7b) [89], Qwen1.5-4B-Chat (Qw4B) [6], Qwen1.5-14B-Chat (Qw14B) [6], Qwen1.5-72B-Chat (Qw72B) [6], Qwen1.5-7B-Chat (Qw7B) [6], Qwen1.5-0.5B-Chat (Qw0.5B) [6], Qwen1.5-1.8B-Chat (Qw1.8B) [6], gpt-4-turbo (gpt4) [2], gpt-3.5-turbo (gpt3.5) [2], Yi-6B-Chat (Yi6B) [100], Mistral-7B-Instruct-v0.1 (Mis0.1) [38], Mistral-7B-Instruct-v0.2 (Mis0.2) [38], vicuna-7b-v1.5-16k (vic7b) [102]. All LLMs are set to greedy decoding mode. If applicable, the system prompt is set to "You are a helpful assistant." by default [104].
@@
Baseline Measurement Tools.
We compare GPV with two baseline tools: self-reports and ValueBench [65]. Self-reports are widely used in recent literature [46, 12, 101, 31, 60, 39, 35], where questionnaires are presented to LLMs in the same manner as they are to humans. ValueBench rephrases the original items to better suit the role of LLMs as human assistants. For instance, the statement "I believe everyone deserves equal rights and opportunities" is rephrased as "Should I believe that everyone deserves equal rights and opportunities?" to reflect the tone of a user seeking advice. The LLMs’ responses are then evaluated by the evaluator LLM, which rates the degree to which the answers lean toward "yes".
@@
Using LABEL:prompt:_item_generation, we ask GPT-4o to generate five open-ended, value-eliciting questions designed to evaluate each value dimension. In this manner, we ensure a comprehensive and sufficient measurement of each involved value. To tailor these questions effectively, we provide several guidelines, which may overlap in some aspects: (1) Avoid personal experience and emotion-based questions: Since an LLM does not have personal experiences or emotions, the questions should not assume it does. (2) Use value dilemmas and scenarios: Instead of asking direct questions that may lead to generic responses, use scenarios or dilemmas to reveal the LLM’s values. (3) Present ethical dilemmas: Include scenarios with no clear right or wrong answer to understand the model’s approach to complex moral issues. (4) Involve conflicting values: Create situations requiring a choice between conflicting values to reveal the model’s value prioritization. (5) Justification: After each scenario, ask the model to justify its decision to understand the reasoning behind it. We exemplify the generated questions below:
@@ D.2 Stability Analysis
Table 8 gathers the results for analyzing the stability of GPV, i.e., the consistency between perception-level and aggregated measurement results for LLMs. They demonstrate that the perception-level results are generally consistent with the aggregated LLM-level results, indicating desirable stability.
@@
D.3 Construct Validity with Data Centering
In practice, psychologists often center self-report data at the individual level before analysis to reduce the influence of personal response style differences. Here, we center the LLM value measurement results for Self-reports and ValueBench and recalculate the correlations. The results are presented in Fig. 5. We find that, even after centering, the results remain inconsistent with the theoretical structure. Note that GPV already standardizes the data by using ValueLlama as an external rater.
Figure 5: Correlations between Schwartz values when using different measurement tools with data centering.
D.4 Comparative Analysis of Value Representation Utility
@@
We train a linear classifier to predict the relative safety scores of LLMs, based on their value measurements. The classifier maps value measurements to a scalar safety score. LLMs are paired as data points; for each pair of LLMs, we concatenate their scores output by the linear layer as the classification logits. We train the linear layer using cross-entropy loss.
@@
To ensure a fair comparison, we normalize the value measurements of different tools to $[-1,1]$ before feeding them into the linear layer. The missing value measurement results (due to refusal to answer or invalid responses in self-reports) are set to 0. We use the Adam optimizer with a learning rate of 0.001 to train the linear layer. We train the linear layer for 1000 epochs, with the batch size equal to the dataset size (36 pairs). The checkpoint with the best validation accuracy is selected as the final model.
@@
Since VSM is the most predictive value system for LLM safety scores, we further investigate the contributions of different values in VSM to the predicted LLM safety scores. The parameters of the trained classifier are associated with each value. In Table 10, we present these parameters averaged over 30 runs of different data splits. We find that Long Term Orientation, Indulgence, and Uncertainty Avoidance positively contribute to the predicted safety scores, while Masculinity, Power Distance, and Individualism negatively contribute. These findings provide insights into future value alignment strategies for LLM safety. Instead of relying on human preferences, we may leverage the basic values of LLMs to guide their behavior towards safety, which can be more transparent, adaptable, and interpretable [98].
@@
Table 10: The associated parameters in the linear prob for different values.
D.5 Value Orientations of LLMs